OpenStudy (anonymous):

Find the coordinates of the missing endpoint if F is the midpoint of E(-3,-6) and F(-5,-7)

OpenStudy (anonymous):

Please Help!!

OpenStudy (anonymous):

someone,.. anyone:(

OpenStudy (anonymous):

(-7,-8)

OpenStudy (anonymous):

how did you get that danieljoel?

OpenStudy (anonymous):

well E(-3,-6) and F(-5,-7) F being the midpoint you dont need a fancy equation to solve this one as simply the x ais has gone -2 and y -1. if F is the midpoint then assume G is the end and go a further -2 on the and -1 on the y giving you G(-7,-8)

OpenStudy (anonymous):

can you give me the actual work? I have to show it:(

OpenStudy (anonymous):

if F is midpoint then as follows x3 = (x2-x1)+x2 and y3 = (y2-y1)+y2 x3= (-5--3)+-5 and y3 = (-7--6)+-7 x3=(-2)-5 and y3 = (-1)-7 x3 = -7 y3 = -8 (-7,-8)

OpenStudy (anonymous):

i just made up a formula, i dont know if there is one for it, im sure there would be but it works :)

OpenStudy (anonymous):

Thanks So much!!

OpenStudy (anonymous):

anytime